1. A number consists of two digits whose sum is 10 . If 72 be subtracted from the number , the digits are reversed . find the number ? 2. Two kg of sugar and 1/2 kg of tea cost Rs.32 and 3 kg of sugar and 1 kg of tea cost Rs.59. Find the cost of 1 kg of tea 1kg of sugar ?
Answer by mukhopadhyay(490)

Let the number be 10x+y
x+y=10;
10x+y-72=10y+x
=>9x-9y=72
=>x-y=8;
We need to solve for x and y from the following equations:
x+y=10
x-y=8
Adding above:
2x=18
=>x=9
y=10-x=1
So, the original number is 91